# Training Long-Context, Multi-Turn Software Engineering Agents with Reinforcement Learning

### v1

@@ 4.1 Main Results
Our two-phase procedure yields substantial improvements. Rejection fine-tuning provides an initial performance boost by enhancing the model’s ability to interact correctly with the environment. This is followed by over 100 RL iterations that progressively refine the agent’s policy. The final model achieves $39.0\%$ on the full SWE-bench Verified. Performance trends are shown in Figure 4.
For head-to-head comparisons, we evaluate DeepSeek-V3-0324, Llama-4 Maverick, Qwen3-235B no-thinking, and Qwen3-32B no-thinking within the same environment and tool setup (see Table 1) on both SWE-bench Verified and SWE-rebench. To benchmark our final model against specialized SWE agents, we summarize recent results in Table 2.
Agent Base Model SWE-bench Ver. before SWE-bench Ver. after Teacher Distillation
@@
DeepSWE-32B, Luo et al. Qwen3-32B 23% 42.2% No
SkyRL-Agent-14B-v0, Cao et al. Qwen3-14B 18.0% 21.6% No
SWE-Fixer-72B, Xie et al. Qwen2.5-72B-Base 30.2% Yes
SWE-agent-LM-32B, Yang et al. Qwen2.5-Coder-32B-Instruct 14.3% 40.2% Yes
Skywork-SWE-32B, Zeng et al. Qwen2.5-Coder-32B-Instruct 6.4% 38.0% Yes
SWE-Gym-32B, Pan et al. Qwen2.5-Coder-32B-Instruct 7.0% 20.6% Yes
SWESynInfer-72B, Ma et al. Qwen2.5-72B-Instruct 25.4% 30.2% Yes
R2EGym-Agent-32B, Jain et al. Qwen2.5-Coder 7.0% 34.4% Yes
Table 2: Comparison of specialized SWE agents on SWE-bench Verified. The “before” column shows the Pass@1 score of the base model prior to training, while the “after” column shows the final performance of the fully trained agent.
Figure 4: A detailed performance trend of RL-trained agent over all iterations. Statistics include Pass@1, Pass@10, the number of submit commands and the average number of steps per trajectory. All metrics are computed on Verified-50.
4.2 Infrastructure
The described process is based on a fully synchronous RL training process, meaning that inference and training stages are interleaved: once rollout generation is completed, trajectories are used for training. This setup enables fully on-policy training with no policy lag between sampling and updates. However, as described earlier, we sample 10 trajectories for each problem. This results in 2-8 optimization steps per iteration depending on batch size and the number of trajectories with $\hat{A}^{(i)}\neq 0$.
We describe more details regarding infrastructure in Appendix C.
4.3 Hyperparameters
Inference.
For rollout generation, we run the model with a temperature of 1.0, explicitly disabling all other decoding parameters such as top_p, top_k, min_p, repetition_penalty and others. This ensures unbiased sampling, which is critical for the validity of importance sampling ratios used during training. We demonstrate the dangers of biased sampling procedures later in our findings.
Training.
Training Stage Problems / Iteration Total Problems Batch Size Clip Range
Stage 1 300 7249 128 [0.2, 0.3]
Stage 2 100 2028 256 [0.2, 0.26]
Table 3: Key hyperparameters across the two RL training stages.
For Qwen2.5-72B-instruct models we use YaRN positional encoding with $\textit{factor}=4.0$ to enable 131k context length training and inference.
@@ 4.4 Findings
A commonly adopted practice in dataset preparation, also mentioned in DeepSWE (Luo et al. 2025), is to filter or mask trajectories that exceed the model’s maximum context length. This is often motivated by the desire to reduce reward noise. However, we find this must be applied with caution. Manually crafted heuristics can introduce biases, breaking the assumption that the training data is sampled from the same distribution as the policy being optimized. In our setting, these long trajectories often occur when the agent is stuck in a repetitive loop. By discarding these trajectories, one also discards specific negative examples of this failure mode. As a result, the agent is not penalized for this looping behavior and fails to learn how to break out of such cycles, which can lead to it occurring more frequently during training.
We also observe a more subtle instability related to discrepancies between sampling and training. Midway through training, we upgraded the vLLM (Kwon et al. 2023) runtime, which introduced internal changes to decoding parameters, enabling top_k and min_p by default. While this initially improved evaluation metrics, it caused performance to degrade after 5–10 training iterations. We attribute this to a mismatch between the rollout and training distributions. The DAPO objective relies on the importance sampling ratio $\rho_{t,k}(\theta)$ to weight the advantage term. This is valid only if rollouts are sampled from $\pi_{old}$. Using different decoding parameters amounts to sampling from a modified distribution, $\pi_{\text{rollout}}\neq\pi_{old}$, which makes the ratio an incorrect estimator and leads to biased gradient updates. Once the sampling configuration was restored, stability recovered.
@@ 5 Discussion and Future Work
Our work successfully demonstrates that modern reinforcement learning algorithms, specifically a modified DAPO, can train capable agents for complex, interactive software engineering tasks. This process, however, highlights fundamental challenges in agent-based learning and reveals several key directions for future research:
@@
  • Sparse Rewards and Credit Assignment. A fundamental challenge is that the agent receives only a single binary success signal at the end of a long trajectory. This sparsity makes it difficult to perform credit assignment, that is, to identify which specific actions in a long sequence were crucial for the outcome. Broadcasting a single advantage estimate across thousands of preceding tokens can result in noisy and inefficient policy updates. Several research directions could address this: (i) reward shaping, which involves designing intermediate rewards based on signals like passing a subset of tests or reducing compiler errors; (ii) training an auxiliary critic or value head to provide step-level advantage estimates, enabling more granular updates; and (iii) prefix sampling, where rollouts are initiated from a shared non-empty trajectory prefix to better isolate the impact of later decisions.
@@
  • Uncertainty and Risk-Awareness. The binary, success-based reward objective encourages the agent to submit a patch “at any cost”, which leads it to act confidently even when a solution is unlikely. For real-world deployment, agents must recognize when to abstain. This requires better uncertainty estimation; for instance, by training the model to explicitly output a confidence score or by using the policy’s output entropy as a proxy for uncertainty. Such estimates would enable a precision-recall trade-off, allowing the agent to decide when to halt or to apply more compute for best-of-$n$ selection without an external outcome-supervision model.
Appendix A Related Work
Software engineering agents.
Early systems, notably SWE-agent (Yang et al. 2024), demonstrated that large language models (LLMs) could effectively operate within sandboxed software environments using predefined toolkits (e.g., shell commands, code editors) and be evaluated via automated unit tests (Jimenez et al. 2024). Subsequent frameworks introduced alternative scaffolding and prompting strategies to enhance model interactions, including Agentless (Xia et al. 2024), OpenHands (Wang et al. 2025), and Moatless (Antoniades et al. 2025). Our work closely aligns with the original SWE-agent design, leveraging similar prompting structures and tool configurations, enabling a direct evaluation of reinforcement learning improvements within this established context.
Strategies to improve SWE agents.
Prior efforts to enhance SWE agent performance beyond improving scaffoldings broadly focus on either test-time exploration or model-level improvements through supervised training. Test-time exploration strategies, such as Monte Carlo Tree Search (MCTS) (Antoniades et al. 2025) and guided 1-step lookahead (Zainullina et al. 2025), significantly boost task success by exploring multiple solution trajectories. However, these methods are computationally expensive and sometimes introduce infrastructure complexity due to operations such as environment rollbacks and parallel execution paths.
Alternatively, many efforts have targeted model-level improvements via supervised fine-tuning (SFT) using expert-curated demonstrations. Prominent examples include SWE-smith (Yang et al. 2025), SWE-Fixer (Xie et al. 2025), and Skywork-SWE (Zeng et al. 2025), all of which achieved success by training open-weight models on extensive demonstration data. In contrast, our method relies exclusively on self-generated interaction data obtained through direct RL training. This approach simplifies data collection, eliminates need for strong teacher models and opens way for iterative self-improvement.
Reinforcement learning for coding.
Reinforcement learning (RL) has shown notable success in structured reasoning domains like mathematics, exemplified by DeepSeekMath (Shao et al. 2024) and Seed-Thinking-v1.5 (Seed et al. 2025). In the coding domain specifically, RL applications traditionally focused on single-turn code generation tasks (Dou et al. 2024), neglecting the complexities introduced by multi-turn interactions and extended contexts.
Recent advancements, however, have started addressing these complexities. Concurrent to our work, DeepSWE (Luo et al. 2025) successfully scaled critic-free RL training to a 32B-parameter model (Qwen3-32B), while Sky-RL (Cao et al. 2025) introduced an asynchronous RL pipeline for long-context tasks. Our work contributes to this emerging area by demonstrating a complete, multi-stage methodology to successfully apply RL to a larger 72B-parameter model with a 131k token context.
@@ Appendix B Verified-50 Problems
Appendix C Infrastructure Details
@@
The entire pipeline of distributed agent execution and evaluation is orchestrated at scale using Kubernetes and Tracto AI (TractoAI 2025). The primary compute units are pods, each provisioned with eight H200 GPUs, 32 CPUs, and 960 GiB of CPU RAM.
@@
Within this environment, model training is conducted using an internal framework built on JAX, while inference is accelerated using the vLLM framework (Kwon et al. 2023), version 0.7.4.

### v2

@@ 5.1 Main results
Our two-phase procedure yields substantial improvements. Rejection fine-tuning provides an initial performance boost by enhancing the model’s ability to interact correctly with the environment. This is followed by over 100 RL iterations that progressively refine the agent’s policy. The performance trends in Figure 3 clearly illustrate how the agent’s behavior changes over iterations and stages. In Stage 1, the agent shows consistent improvement before its Pass@1 score begins to plateau. The switch to Stage 2 provides a further performance increase. Notably, this second stage is also characterized by a significant growth in the average steps per trajectory and the number of trajectories finished by a submit command, suggesting that the agent engages in longer, more complex reasoning to solve the more difficult tasks. The final model achieves $39.0\%$ on the full SWE-bench Verified.
For head-to-head comparisons, we evaluate DeepSeek-V3-0324, Llama-4 Maverick, Qwen3-235B-A22B no-think, Qwen3-32B no-think, gpt-oss-120b and Qwen3-235B-A22B-Instruct-2507 within the same environment and tool setup (see Table 1) on both SWE-bench Verified and SWE-rebench. To benchmark our final model against specialized SWE agents, we summarize recent results in Table 2.
Table 2: Comparison of specialized multi-turn SWE agents on SWE-bench Verified. The “Before” column shows the Pass@1 score of the model prior to training, while the “After” column shows the final performance. “Teacher Distillation” indicates whether a stronger model was used for training.
Agent Base Model Before After Teacher Distillation
@@
DeepSWE-32B, (Luo et al., 2025) Qwen3-32B 23.0% 42.2% No
SkyRL-Agent-14B-v0, (Cao et al., 2025) Qwen3-14B 18.0% 21.6% No
SWE-Fixer-72B, (Xie et al., 2025) Qwen2.5-72B-Base 30.2% Yes
SWE-agent-LM-32B, (Yang et al., 2025) Qwen2.5-Coder-32B-Instruct 14.3% 40.2% Yes
Skywork-SWE-32B, (Zeng et al., 2025) Qwen2.5-Coder-32B-Instruct 6.4% 38.0% Yes
SWE-Gym-32B, (Pan et al., 2025) Qwen2.5-Coder-32B-Instruct 7.0% 20.6% Yes
SWESynInfer-72B, (Ma et al., 2024) Qwen2.5-72B-Instruct 25.4% 30.2% Yes
R2EGym-Agent-32B, (Jain et al., 2025) Qwen2.5-Coder-32B-Instruct 7.0% 34.4% Yes
@@ 5.2 Findings
A commonly adopted practice in dataset preparation, also mentioned in DeepSWE (Luo et al., 2025), is to filter or mask trajectories that exceed the model’s maximum context length. This is often motivated by the desire to reduce reward noise. However, we find this must be applied with caution. Manually crafted heuristics can introduce biases, breaking the assumption that the training data is sampled from the same distribution as the policy being optimized. In our setting, these long trajectories often occur when the agent is stuck in a repetitive loop. By discarding these trajectories, one also discards specific negative examples of this failure mode. As a result, the agent is not penalized for this looping behavior and fails to learn how to break out of such cycles, which can lead to it occurring more frequently during training.
We also observe a more subtle instability related to discrepancies between sampling and training. Midway through training, we upgraded the vLLM (Kwon et al., 2023) runtime, which introduced internal changes to decoding parameters. The upgrade enabled top_k and min_p filtering (previously off by default) using model-dependent values inherited from Hugging Face configurations. While this initially improved evaluation metrics, it caused performance to degrade after 5–10 training iterations. We attribute this to a distribution mismatch that violates the assumptions of importance sampling. The DAPO objective relies on the probability ratio $\rho_{t,k}(\theta)$ to correct for evaluating actions under the current policy $\pi_{\theta}$ using data generated by the policy from the previous iteration $\pi_{\theta_{old}}$ (see Equation 6). Enabling decoding filters like top_k or min_p means that trajectories were sampled from a modified, truncated distribution $\pi_{\text{rollout}}$, not the true policy $\pi_{\theta_{old}}$. Consequently, the ratio $\rho_{t,k}(\theta)$ becomes an invalid estimator of the policy change, leading to biased gradient updates and training instability. Once unbiased sampling was restored, performance recovered.
@@ 6 Discussion and future work
Our work successfully demonstrates that modern reinforcement learning algorithms, specifically those based on the DAPO framework, can train capable agents for complex, interactive software engineering tasks. This process, however, highlights fundamental challenges in agent-based learning and reveals several key directions for future research:
@@
  • Sparse Rewards and Credit Assignment: A fundamental challenge is that the agent receives only a single binary success signal at the end of a long trajectory. This sparsity makes it difficult to perform credit assignment, that is, to identify which specific actions in a long sequence were crucial for the outcome. Broadcasting a single advantage estimate across thousands of preceding tokens can result in noisy and inefficient policy updates. Several research directions could address this: (i) reward shaping, which involves designing intermediate rewards based on signals like passing a subset of tests or reducing compiler errors; (ii) training an auxiliary critic or value head to provide step-level advantage estimates, enabling more granular updates; and (iii) prefix sampling, where rollouts are initiated from a shared non-empty trajectory prefix to better isolate the impact of later decisions.
@@
  • Uncertainty and Risk-Awareness: The binary, success-based reward objective encourages the agent to submit a patch “at any cost”, which leads it to act confidently even when a solution is unlikely. For real-world deployment, agents must recognize when to abstain. This requires better uncertainty estimation; for instance, by training the model to explicitly output a confidence score or by using the policy’s output entropy as a proxy for uncertainty. Such estimates would enable a precision-recall trade-off, allowing the agent to decide when to halt or to apply more compute for best-of-$n$ selection without an external outcome-supervision model.
Appendix A PPO, GRPO and DAPO objectives
In our setting, the PPO objective can be the following:
$\mathcal{J}_{\text{PPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}o_{0}\sim\mathcal{D},\\ \tau\sim\pi_{\theta_{old}}(\cdot|o_{0})\end{subarray}}\Biggl[\frac{1}{|\tau|}\sum_{t=0}^{|\tau|-1}\frac{1}{|a_{t}|}\sum_{k=1}^{|a_{t}|}\min\Bigl(\rho_{t,k}(\theta)A_{t,k},\text{clip}(\rho_{t,k}(\theta),1-\varepsilon,1+\varepsilon)A_{t,k}\Bigr)\Biggr],$ (5)
where $o_{0}\sim\mathcal{D}$ indicates that initial observations are sampled from the distribution of tasks in our dataset and $\theta_{old}$ represents parameters of the policy from the previous training step.
$\rho_{t,k}(\theta)\triangleq\frac{\pi_{\theta}(a_{t,k}|h_{t},a_{t,<k})}{\pi_{\theta_{\text{old}}}(a_{t,k}|h_{t},a_{t,<k})}$ (6)
is the probability ratio for the $k$-th token in action $a_{t}$, and $A_{t,k}$ is its corresponding advantage estimate.
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}o_{0}\sim\mathcal{D},\\ \{\tau^{(i)}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|o_{0})\end{subarray}}\Biggl[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|\tau^{(i)}|}\sum_{t=0}^{|\tau^{(i)}|-1}\frac{1}{|a_{t}^{(i)}|}\sum_{k=1}^{|a_{t}^{(i)}|}\\ \min\Bigl(\rho_{t,k}^{(i)}(\theta)\hat{A}^{(i)},\text{clip}(\rho_{t,k}^{(i)}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}^{(i)}\Bigr)\Biggr].$ (7)
DAPO suggests modifications on top of GRPO described in Section 3.2:
$\mathcal{J}_{\text{DAPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}o_{0}\sim\mathcal{D},\\ \{\tau^{(i)}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|o_{0})\end{subarray}}\Biggl[\frac{1}{\sum_{i=1}^{G}\sum_{t=0}^{|\tau^{(i)}|-1}|a_{t}^{(i)}|}\sum_{i=1}^{G}\sum_{t=0}^{|\tau^{(i)}|-1}\sum_{k=1}^{|a_{t}^{(i)}|}\\ \min\Bigl(\rho_{t,k}^{(i)}(\theta)\hat{A}^{(i)},\text{clip}(\rho_{t,k}^{(i)}(\theta),1-\varepsilon_{low},1+\varepsilon_{high})\hat{A}^{(i)}\Bigr)\Biggr].$ (8)
Figure 3: A detailed performance trend of the RL-trained agent over all iterations. Statistics include Pass@1, Pass@10, the number of submit commands and the average number of steps per trajectory. All metrics are computed on Verified-50.
@@ Appendix B Verified-50 problems
Listing 1: A list of problems from Verified-50.
Appendix C Hyperparameters
Inference.
For rollout generation, we run the model with a temperature of 1.0, explicitly disabling all other decoding parameters such as top_p, top_k, min_p, repetition_penalty and others. This ensures unbiased sampling, which is critical for the validity of importance sampling ratios used during training. We demonstrate the dangers of biased sampling procedures in Section 5.2.
Training.
Table 3: Key hyperparameters across the two RL training stages.
Hyperparameter Stage 1 Stage 2
Problems / Iteration 300 100
Total Problems 7249 2028
Batch Size 128 256
($\varepsilon_{low}$, $\varepsilon_{high}$) (0.2, 0.3) (0.2, 0.26)
Maximum Turns ($T_{\text{max}}$) 40 80
Penalty Threshold ($L_{\text{thr}}$) 10 10
For the Qwen2.5-72B-Instruct model we use YaRN positional encoding (Peng et al., 2024) with $\textit{factor}=4.0$ to enable 131k context length training and inference.
Appendix D Infrastructure details
The described process is based on a fully synchronous RL training process, meaning that inference and training stages are interleaved: once rollout generation is completed, trajectories are used for training. This setup enables fully on-policy training with no policy lag between sampling and updates. However, as described earlier, we sample 10 trajectories for each problem. This results in 2-8 optimization steps per iteration depending on the batch size and the number of trajectories with $\hat{A}^{(i)}\neq 0$.
@@ Appendix D Infrastructure details
The entire pipeline of distributed agent execution and evaluation is orchestrated at scale using Kubernetes for agent execution and Tracto AI (TractoAI, 2025) for evaluation. During the rollout generation phase, each agent instance runs in a dedicated Kubernetes pod with a resource request of 0.5 CPU and 2 GiB of RAM.
All training and inference are conducted on a cluster of 16 H200 nodes, each equipped with eight GPUs, 32 CPUs and 960 GiB of CPU RAM.
@@
Within this environment, model training is conducted using an internal framework built on JAX, while inference is accelerated using the vLLM framework (Kwon et al., 2023), version 0.7.4.